OpenStudy (anonymous):

x+y=136 x=3y find the intercepts and slop for each

OpenStudy (amistre64):

slope of the top is: -1 ; yint = 136 slope for the bottom is: 1/3 ;yint = 0

OpenStudy (anonymous):

I am sorry i am so lost when it comes to this. When it states to find the intercepts does it mean only the (Y)?

OpenStudy (amistre64):

the y intercept is where the graph cross over and gets intercepted by the y axis..... the y axis is also know as x=0 so wherever x=0, the rest of the equation has to equal out. for example: x = 3y 0 = 3y....yint = 0 x+y = 136 0 + y = 136 .... yint = 136

OpenStudy (anonymous):

so it would be 3,0 and 0,136? The reason why i am asking is because the question that follows is, if you graph these two will they intersect and if so where?

OpenStudy (amistre64):

close, real close..but remember that the first coordinate is your x value; and since yint means x=0 we get: (0,0) and (0,136)

OpenStudy (anonymous):

So then the question asks for the intercepts i would give just 136 for the first one and 0 for the next one or is the proper way to right that (0,136) and (0,0)?

OpenStudy (anonymous):

write***

OpenStudy (amistre64):

i think the proper way is with the "coordinate" notation: (0,0) (0,136) but it just depends on how the "test" wants you to write them....

OpenStudy (anonymous):

ok so when it asks if they intersects i would say yes, and they intersect at 0?